A 1.00 liter solution contains 0.34 M ammonia and 0.44 M ammonium bromide.

If 0.110 moles of barium hydroxide are added to this system, indicate whether the following statements are true or false.

(Assume that the volume does not change upon the addition of barium hydroxide.)

_______TrueFalse A. The number of moles of NH3 will increase.

_______TrueFalse B. The number of moles of NH4+ will remain the same.

_______TrueFalse C. The equilibrium concentration of H3O+ will remain the same.

_______TrueFalse D. The pH will decrease.

_______TrueFalse E. The ratio of [NH3] / [NH4+] will increase.

A 1.00 liter solution contains 0.52 M ammonia and 0.40 M ammonium iodide.

If 0.2 moles of hydrobromic acid are added to this system,
indicate whether the following statements are true or false.

(Assume that the volume does not change upon the addition of hydrobromic acid.)

_______TrueFalse A. The number of moles of NH3 will decrease.

_______TrueFalse B. The number of moles of NH4+ will decrease.

_______TrueFalse C. The equilibrium concentration of H3O+ will increase.

_______TrueFalse D. The pH will remain the same.

_______TrueFalse E. The ratio of [NH3] / [NH4+] will remain the same.

Submit Answer

Retry Entire Group

A 1.00 liter solution contains 0.49 M hydrofluoric acid and 0.37 M potassium fluoride.

If 25.0 mL of water are added to this system, indicate whether the following statements are true or false.

(Note the the volume MUST CHANGE upon the addition of water.)

_______TrueFalseA. The concentration of HF will remain the same.

_______TrueFalseB. The concentration of F- will increase.

_______TrueFalseC. The equilibrium concentration of H3O+ will decrease.

_______TrueFalseD. The pH will decrease.

_______TrueFalseE. The ratio of [HF] / [F-] will remain the same.
Answer #1

1) True A. The number of moles of NH3 will increase.

False B. The number of moles of NH4+ will remain the same.

False C. The equilibrium concentration of H3O+ will remain the same.

False D. The pH will decrease.

True E. The ratio of [NH3] / [NH4+] will increase.

Before addition of Ba(OH)2

pH of basic buffer = pka+ log(NH3/NH4Br)

pka of NH3 = 14-4.75 = 9.25

no of mol of NH4Br = 0.44*1 = 0.44 mol

no of mol of NH3 = 0.34*1 = 0.34 mol

pH = 9.25+log(0.34/0.44)

pH = 9.14

after addition of Ba(OH)2

pH of basic buffer = pka+ log(NH3 + OH-/NH4Br - OH-)

pka of NH3 = 14-4.75 = 9.25

no of mol of NH4Br = 0.44*1 = 0.44 mol

no of mol of NH3 = 0.34*1 = 0.34 mol

no of mol of OH- added = 0.11*2 = 0.22 mol ( 1mol Ba(OH)2 = 2 mole OH-)

pH = 9.25 + log((0.34+0.22)/(0.44-0.22))

pH = 9.66

as pH increases,H3O+ Concentration decreases.and also he ratio of [NH3] / [NH4+] will increase.

2)

True A. The number of moles of NH3 will decrease.

False B. The number of moles of NH4+ will decrease.

True C. The equilibrium concentration of H3O+ will increase.

False D. The pH will remain the same.

False E. The ratio of [NH3] / [NH4+] will remain the same.

Before addition of HBr

pH of basic buffer = pka+ log(NH3/NH4Br)

pka of NH3 = 14-4.75 = 9.25

no of mol of NH4I = 0.4*1 = 0.4 mol

no of mol of NH3 = 0.52*1 = 0.52 mol

pH = 9.25+log(0.52/0.4)

pH = 9.36

after addition of HBr

pH of basic buffer = pka+ log(NH3 - H+/NH4Br + H+)

pka of NH3 = 14-4.75 = 9.25

no of mol of NH4Br = 0.44*1 = 0.4 mol

no of mol of NH3 = 0.34*1 = 0.52 mol

no of mol of HBr added = 0.2 mol

pH = 9.25 + log((0.52-0.2)/(0.4+0.2))

pH = 8.98

as pH decreases,H3O+ Concentration increases.and also he ratio of [NH3] / [NH4+] will decrease.

3)

False A. The concentration of HF will remain the same.

False B. The concentration of F- will increase.

False C. The equilibrium concentration of H3O+ will decrease.

False D. The pH will decrease.

True E. The ratio of [HF] / [F-] will remain the same.

pH of acidic buffer = pka + log([KF]/[HF])

Addition of water changes concentration of KF,HF but the ratio of [KF]/[HF] will remains same.so that, there is no change in pH, H3O+ concentration